OpenStudy (anonymous):

How would you solve 1/cot2x+1?

OpenStudy (anonymous):

cot2x = cos2x/sin2x Then cos2x = 2cosxsinx I think.....similarly sin2x......So, you can keep going at it until it is rather simplified nicely.

OpenStudy (anonymous):

i don't understand how you got cos2x = 2cosxsinx... @GT

OpenStudy (anonymous):

It was one of the theorem thingies....I don't remember....was just getting you started on the path. Need to look up on google to confirm.

OpenStudy (anonymous):

don't see that theorem anywhere in our identites box in our textbook.

OpenStudy (anonymous):

My mistake. Here is the right one. cos(2x) = cos(x+x) = cos(x)cos(x) - sin(x)sin(x) = cos²(x) - sin²(x)

OpenStudy (anonymous):

the property is cos2x + sin2x = 1 though....?

OpenStudy (anonymous):

@saifoo.khan can you please help with this one? I am wandering off into some weeds that don't get to the answer.

OpenStudy (anonymous):

cos^2x + sin^2x = 1. Correct. That is "squared".

OpenStudy (anonymous):

beyond confused, thanks for baring with me through all the confusion..

OpenStudy (anonymous):

Is that what you meant by cot2x?

OpenStudy (anonymous):

i meant it cot^2x, forgot that sign!

OpenStudy (anonymous):

cot^2x is not same as cot2x. When I saw cot2x, I thought it was cot(2x).

OpenStudy (anonymous):

so sorry! typo!

OpenStudy (anonymous):